Find the coordinate of those points which lie on the parabola y^2=8x and whose focal distance is 4

Grade:12

1 Answers

Nimish bansal
6 years ago
Focal distance = x1 + a. Where a is the x abcissa of focus and x1 is the abcissa of point on parabola from which we are measuring the focal distance. 4= x1 + 2x1= 2y1^2=8×2y1=+/- 4Point(+/-4,2)